DX LUCI AO Commissioning

00:58
  • Waiting for the sky to get dark enough.
01:19
  • Checking pointing.
  • Cancelled preset.
  • Couldn't find the star.
01:25
  • New preset, the system has found the star.
  • Bin 3, 400 Hz.
  • 12.9 mag in the preset
  • 13.5 in the ccd 39
  • Optimize Gain. 1.94 2.29 1.10
  • Manual gain: 1.94 1.19 1.00
  • No GOPT and NCPA Correction.
01:51
  • Running stably, waiting for a LUCI problem
02:46
  • Alexander fixed the LUCI problem.
  • Datasets: 20171102_030016, 030033, 030055, 030108
04:06
  • New preset
  • mag 14.4 in the preset.
  • 15.1 bin 4, 200Hz
  • Optimized Gain: 3.40 2.00 3.00
  • Manual gains: 3.40 1.20 1.10
  • The Camera lens is moving to the wrong place.
  • I added to the IT #6856.
05:34
  • Still working.
05:55
  • SX had a skip frame.
  • We paused the script before the loop failed.
06:04
  • Getting a few % safe skip.
  • There are current warnings for actuator 338
  • The seeing was over 1", now it's 0.9"
  • No safe skips but still warnings about 338
06:07
  • Another few percent safeskip
06:11
  • manually adjusted the pupil position, to see if it helped, it didn't seem to.
06:22
  • Diled back the TT gain to 2.40, that seems to have reduced the frequency of safeskips.
  • There are still several when the seeing gets above 1.1"
06:27
  • Loop opened due to safe skips. I re-closed with lower gains. 1.70 0.90 0.90
  • It's much more stable.
07:13
  • Finished up

-- %USERSIG{GregoryTaylor - 2017-11-02}%
Topic revision: r2 - 02 Nov 2017, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback